A reaction vessel contains 0.0200 M N2, 0.0600 M H2, and 0.0100 M NH3. With the aid of calculations, predict whether ammonia, NH3, will be formed or dissociate when the mixture goes to equilibrium at 400°C. Kc for the reaction is 0.500 at 400°C. Reaction: N2 (g) + 3H2 (g) 2NH3 (g)


1
Expert's answer
2021-10-13T04:47:34-0400

0.06*400/0.5*400=24/200=0.12